How would I know if the assumptions for this test have been met?
A new brand of car gets 35 miles per gallon on average. Suppose a consumer group claims that these cars get less than 35 miles per gallon. A sample of 40 cars is selected, and the sample mean for the 40 cars is 33 miles per gallon while the sample standard deviation is 3.8

Answer by ewatrrr(24785)

How would I know if the assumptions for this test have been met? n > 30
H0: mu = 35m/g
H1: mu < 35 m/g
So we're conducting a one tailed test. The p value will visually represent the area under the curve in the left tail (that's to the left of the test statistic)
-------------------------------------------------------
Now compute the test statistic:
z = (x - mu)/(s/sqrt(n))
z = (33 - 35)/(3.8/sqrt(40)) = -3.3287
Finding the p-value invNorm(-3.3287) = .0004
You did not state a confidence level : .05 commonly used
.0004 < .05, reject the null htpothesis